What is your predicted distribution of genotypes for a total of 12387, assuming the population is at hardy-weinberg equilibrium? Given p=0.877 and q=0.123

The answer to this question is A/A = 9527, A/B = 2672 and B/B = 187. Could you explain how to derive this answer.

Thank you!!
Answer #1

Let, p= frequency of dominant allele

q= frequency of recessive allele

p2 = frequency of homozygous dominant individual

2pq= frequency of heterozygous individual

q2 = frequency of homozygous recessive individual

According to Hardy- Weinberg rules,

p+q=1 and p2 + 2pq+ q2 = 1.

Given, p=0.877 and q=0.123

p2 = 0.877x 0.877=0.7691.

Total is 12387 so A/A (p2)= 12387x 0.7691= 9527

q2 = 0.123x 0.123= 0.015129

Total is 12387 so B/B (q2 )= 12387x 0.015129= 187

2pq= 2x 0.877x 0.123= 0.215742

Total is 12387 so A/B (2pq) = 12387x 0.215742= 2672.